OpenStudy (anonymous):

What is the length of the midsegment of the trapezoid made by the vertices A(0, 5), B(3, 3), C(5, -2) and D(-1, 2)

563blackghost (563blackghost):

First we would find the length of AB so we use distance formula.... \(\LARGE{d=\sqrt{(3-0)^{2}+(3-5)^{2}}}\)

563blackghost (563blackghost):

And to find the length of CD we do the same equation.... \(\LARGE{\sqrt{d=(-1 - 5)^{2}+(2-(-2))^{2}}}\) Once you find the length of AB and CD you would then add the two lengths together and then divide it by 2.... \(\Huge{Midsegment=\frac{AB+CD}{2}}\)